OpenStudy (anonymous):

Write each expression as the product of two binomials x2 - 5x + 6.. how do I do it ? not just the answer..

OpenStudy (anonymous):

(x-3)(x-2)

OpenStudy (anonymous):

Not just the answer pleasee I really need to learn this ..

OpenStudy (diyadiya):

x^2 -3x-2x+6=0 x(x-3)-2(x-3)=0 (x-2)(x-3)=0

OpenStudy (diyadiya):

oh no need of that zero as you are not finding the value of x

OpenStudy (anonymous):

How did you get the last one the (x - 2)(x - 3) ?

OpenStudy (anonymous):

and how do you know when its positive and negative?

OpenStudy (diyadiya):

x-3 we take common and rest is x-2 :)

OpenStudy (anonymous):

its pretty much finding the two numbers whos product equals 6 abnd whose sum is -5

OpenStudy (anonymous):

Thank youu so muchh guys :)

OpenStudy (diyadiya):

coefficient of 'x' is sum of two numbers i.e -5 here and '+6' is the product part so accordingly we put the -ve and +ve signs

OpenStudy (anonymous):

Ohhh I was just looking at it and I kept trying to figure out where the 5 had went but it took from the 6..

OpenStudy (anonymous):

Ohh noo.. actually you added a 2x that wasnt supposed to be on there..

OpenStudy (diyadiya):

we just split the middle term

OpenStudy (diyadiya):

try remembering it like this x^2-Sx+P , where S = sum of two numbers and P= Product of two numbers ..

OpenStudy (diyadiya):

sorry x^2+Sx+P

OpenStudy (anonymous):

wait what are you trying to figure out

OpenStudy (anonymous):

Uff I feel dumb it all looks chinese and im trying to Write each expression as the product of two binomials x2 - 5x + 6

OpenStudy (anonymous):

lol nah i was a little confused with what diya said too lmao

OpenStudy (diyadiya):

oh sorry if i made you confused

OpenStudy (anonymous):

Nooo its okayy im easily confused.. its not your fault at alll your just trying to help me so thank youu :)

OpenStudy (anonymous):

factoring is hard to explain its like a puzzle finding which numbers fit the best

OpenStudy (diyadiya):

welcome.. if u want i'll TRY to explain it to u again :) this is the way we learnt ..

OpenStudy (anonymous):

Lol yes it looks hard to explain and yes pleasee :)

OpenStudy (anonymous):

the first thing is finding numbers that multiply to form 6 examples: 2 times 3, 1 times 6

OpenStudy (anonymous):

but the problem is you have to arrange one of those pairs of numbers so it add up to to equal the middle term -5

OpenStudy (anonymous):

Yes its easy for me to find the numbers I just dont know the order of where the numbers go

OpenStudy (diyadiya):

first try to factor the product part for example here its 6 factors of 6 are ,1,2,3,6 then take two suitable numbers which will give you the sum

OpenStudy (anonymous):

-1 x 6 ? would that equal -5 right

OpenStudy (diyadiya):

no -1 * 6 = -6

OpenStudy (anonymous):

I know multiplied equals 6 but added it equals 5 ?

OpenStudy (anonymous):

the middle term..

OpenStudy (anonymous):

so you get -2 and -3 though right is that what your saying not sure

OpenStudy (diyadiya):

we have to see the product too if u take -1 and 6 it will give you product ' -6 ' and sum 5 ...we need -5 as sum and +6 as product

OpenStudy (anonymous):

Ohhh okayy I understand :) after we get the numbers where do they go ?

OpenStudy (diyadiya):

we write those numbers in the middle .. -2x + -3x = -5x

OpenStudy (diyadiya):

=x^2 -3x-2x+6

OpenStudy (anonymous):

Ohhh I see how you got the first part of the equation now

OpenStudy (diyadiya):

x(x-3)-2(x-3) this part ?

OpenStudy (anonymous):

noo the one you just said lol.. how do you get that part ?

OpenStudy (anonymous):

lol too much to read lmao sorry guys im doing my math hw in btw lol

OpenStudy (diyadiya):

u split the middle term that is -5 with the numbers u just got by factorising the product part i.e 6

OpenStudy (diyadiya):

* middle term is" -5x "

OpenStudy (anonymous):

lala if you want email me: pyrrolysinesynthesis@gmail.com

OpenStudy (diyadiya):

so we write it as sum of those numbers u got .. -5x = -2x + -3x got it ?

OpenStudy (anonymous):

Yes..

OpenStudy (diyadiya):

if u didn't then tell me ok :) so u get the equation x^2 -3x-2x+6

OpenStudy (anonymous):

yes I understood that but I thought you were explaining the next part so I was trying to figure out how it applied .. Lol.. but I got everything so far :)

OpenStudy (diyadiya):

ok x^2 -3x-2x+6 next part.. we 1st consider first two terms x^2 -3x and take the common out ..so we get " x(x-3) then take the last two terms -2x + 6 take the common out ..so we get -2(x-3) got it ? :D

OpenStudy (diyadiya):

and we write it together

OpenStudy (diyadiya):

as x(x-3)-2(x-3

OpenStudy (diyadiya):

??

OpenStudy (anonymous):

sorry my comp stopped working for a second

OpenStudy (anonymous):

but yes I understand how you got thatt

OpenStudy (anonymous):

Thank youu soo muchh by the way :D If I could give you 78592 medals I would Lol. i understand the proccess now.. now I just have to practice it on the other problems I have for hw

OpenStudy (diyadiya):

oh you are most welcome :D lala All the Best :)